OpenStudy (anonymous):

Two boys want to use a seesaw and they need to move the seesaw so that their weights will balance out. The formula is given by w1 • d1 = w2 • d2 where w1 = weight of the first boy, d1 = distance of the first boy from the fulcrum, w2 = weight of the second boy, and d2 = distance of the second boy from the fulcrum. Rewrite the formula to solve for d2. Show all steps in your work.

OpenStudy (anonymous):

Well since you know the formula of w1 * d1 = w2 * d2, can't you just divide all of it by w2 to get your answer? (w1 * d1)/w2 =d2

OpenStudy (anonymous):

um maybe lol I have no idea how to do any of this :(

OpenStudy (radar):

Sounds reasonable to me @darthspectrum